# Large Language Models Hallucination: A Comprehensive Survey

### 09 Oct 2025 v2

@@ 6.2 Retrieval-Based Techniques
Retrieving external knowledge for LLMs can enhance the factual grounding, contextual relevance, and overall accuracy of LLMs. The information contained in the model parameters during pre-training is known as parametric knowledge. This knowledge is static and may not encompass the most recent or domain-specific information. To overcome this limitation, non-parametric knowledge can be incorporated through retrieval-based techniques. In the context of LLMs, knowledge retrieval enables LLMs to dynamically access external verified sources of data, search engines, databases, and specific text corpora. By interacting with these reliable sources, models can consult up-to-date and domain-specific data during generation, which in turn improves the factual accuracy of their outputs [130, 131]. Retrieval-based techniques can be classified into three categories based on the phase of accessing external resources: before, during, and after generation [9]. We used these knowledge retrieval stages to categorize retrieval-based techniques proposed for hallucination mitigation.
@@ 6.2.2 Knowledge Graph
Inference stage. KGs can also mitigate hallucinations in the inference stage by employing KG-augmented retrieval. This allows LLMs to retrieve the appropriate subgraph for answering a query grounded by real-world facts. KAPING [150] retrieves related triples from KGs for zero-shot QA by matching entities in questions. Similarly, StructGPT [151] enhances LLM responses using data from KGs, tables, and databases and employing structured queries for information retrieval. For complex reasoning tasks, IRCoT [152] integrates KGs with CoT to iteratively guide retrieval and reasoning for multi-step questions. Likewise, RoG [153] uses KGs to construct reliable reasoning pathways grounded in diverse relationships to improve accuracy. For real-time applications, FactGenius [154] combines LLM-based connection filtering with Levenshtein distance validation, which improved fact verification F1-scores by 12% on the FactKG benchmark through fuzzy relation mining. Moreover, KG-based retrofitting (KGR) [155] incorporates LLMs with KGs to mitigate factual hallucination during the reasoning process. Unlike previous methods that only use the user’s input to query the knowledge graph, KGR retrofits the initial draft responses of LLMs based on the factual knowledge stored in KGs. The framework leverages LLMs to extract, select, validate, and retrofit factual statements within the model-generated responses, which enables an autonomous knowledge-verification and refining procedure without additional manual efforts.
@@ 6.3 Reasoning-Based Techniques
CoT Reasoning. CoT is an improved version of few-shot prompting [121] that follows a step-by-step approach in natural language to decompose multi-step problems into intermediate steps. This method provides an interpretable view of the model’s reasoning process, which enables better insight into how the model arrives at its conclusions. Zhao et al. [156] proposed a Zero-shot-CoT reasoning method by incorporating logical thoughts (LoT), which employs symbolic logic to validate and revise each step. Therefore, it directly addresses and mitigates hallucinations arising from logical flaws. Sultan et al. [157] introduced Structured CoT (SCoT), which utilizes a structured state-machine methodology to systematically oversee content reading, utterance generation, and hallucination mitigation phases. This method improves the model’s faithfulness and reduces hallucinations in content-grounded tasks. Li et al. [158] further adapted the SCoT approach specifically for code generation tasks, introducing programming structures, such as sequential, branching, and looping steps, as intermediate representations. Their approach explicitly guides the LLMs to follow structured programming logic, which enhances the logical correctness, clarity, and readability of the generated code. This structured reasoning substantially mitigates hallucinations and improves code accuracy, outperforming traditional CoT prompting by a considerable margin across multiple programming benchmarks.
@@
Self-Consistency. Building on the success of CoT, Wang et al. [159] introduced a self-consistency approach, which replaces the decoding strategy of CoT. Unlike the original CoT method, the authors proposed the concept of generating multiple reasoning chains for a certain topic. The final answer is determined by conducting a majority vote or by selecting the most consistent response from several chains. Li et al. [160] extended self-consistency through proposing DIVERSE. This method generates multiple reasoning pathways through diverse prompts, systematically validates each reasoning step, and utilizes weighted voting according to step consistency. The step-aware approach significantly reduces hallucinations by isolating and rectifying errors at the step level rather than the whole chain. Similarly, Liang et al. [161] proposed an RL from knowledge feedback (RLKF) technique, which integrates self-consistency and internal knowledge state assessments. RLKF enhances the model’s ability to resist factual hallucinations by reinforcing the consistency between the internal knowledge state and external outputs. Xu et al. [162] proposed SaySelf, a method for training LLMs to produce self-reflective rationales conditioned on inconsistencies across multiple sampled reasoning trajectories. SaySelf explicitly prompts the model to provide precise confidence estimates and rationales that detail its knowledge gaps and uncertainty. This method substantially enhances confidence calibration and effectively mitigates hallucinations.
@@
Iterative Refinement. CoT has substantially improved the reasoning abilities of LLMs, hence reducing hallucination. Nevertheless, if CoT initiates the sequence with erroneous reasoning, it fails to rectify reasoning errors or factual inaccuracies that may occur during the reasoning process. This limitation often leads to hallucinations. Inspired by the way people edit their written content, Madaan et al. [122] proposed Self-Refine to enhance the generated output by an LLM through iterative refinement, which depends on domain-specific data, external supervision, and RL. However, these techniques require large annotated datasets that are unavailable for several domains. Self-Refine leverages the concept of LLMs as self-reviewers by using a single LLM that acts as a generator, refiner, and feedback provider, relying solely on prompt designs without the need for fine-tuning. Self-Refine depends on an appropriate LLM and three prompts (for initial response generation, feedback, and refinement). It works by generating an initial output based on an input sequence, providing feedback on the output, then subsequently refining the output in accordance with the self-generated feedback. It alternates between feedback and refining until a specified condition is achieved. Similarly, Shinn et al. [163] proposed Reflexion method that involves finding the subsequent optimal solution in planning using ReAct [164]. The Reflexion uses a long-lasting memory that allows an agent to recognize its own mistakes and autonomously derive insights from its mistakes and iteratively adapt its behavior over time. Instead of using RL, Reflexion leverages an external verbal feedback mechanism. This technique operates by employing an actor that generates an output and an evaluator to evaluate the generated output. The process continues until the evaluator thinks the agent’s output is correct or when a maximum number of trials is reached. In contrast to Self-Refine, Reflexion explicitly integrates external evaluations and memory mechanisms, which enhances learning effectiveness over both short-term and long-term experiences. Paul et al. [165] developed REFINER by applying structured intermediate criticism from a critic model. The critic model critiques single reasoning steps and allows iterative refinement by identifying and correcting inaccuracies throughout the reasoning process. Likewise, Lee et al. [166] introduced the Ask, Assess, and Refine (A2R) methodology, which methodically assesses outputs for factual accuracy and hallucinations through metric-based feedback. A2R employs natural language feedback from metric assessments, which progressively enhances responses according to metrics like accuracy, fluency, and citation precision to reduce hallucination. In another study, Zhang et al. [167] introduced PREFER, a refinement-focused ensemble method, which works by continuously merging and refining prompts through identifying the limitations of previous iterations, enhancing model stability and overall output quality. In addition, Huang et al. [168] demonstrated that LLMs can enhance their reasoning capabilities by generating high-confidence rationale-augmented responses. Their approach integrates unsupervised iterative refinement with self-consistency and feedback-based fine-tuning, which significantly improved reasoning performance.
@@ 6.4 Model-centric Training and Adaptation
Decoding Strategies. Decoding is the mechanism used to convert encoded representations of the LMs’ output into comprehensible text. During decoding, the model iteratively selects tokens from its vocabulary, which builds contextually relevant and syntactically accurate sentences. However, decoding strategies may contribute to generating hallucinated text [73]. Selecting an effective decoding technique can help the model to generate outputs that are more grounded in context and aligned with user expectations. Recent studies have shown promising results in mitigating hallucinations by building upon the on-the-shelf decoding strategies. Chen et al. [175] introduced DoLa to mitigate hallucination in LLMs without the need for fine-tuning or external retrieval mechanisms. It enhances factual accuracy by leveraging differences between mature (higher) and premature (lower) transformer layers. Evaluated on the TruthfulQA benchmark, DoLa attained a 12–17% enhancement in truthfulness scores across various LLMs. Similarly, Shi et al. [176] enhanced the contextual faithfulness of text generation by reducing the influence of prior knowledge during the decoding process. They proposed a contrastive decoding technique that modifies output probabilities to enhance the influence of the given context. The context-aware decoding operates by contrasting the output probabilities of the model when the context is included and excluded from the prompt. This ensures that the model gives more weight to contextually relevant tokens, reducing reliance on outdated or incorrect prior knowledge. This approach has been evaluated on a summarization task, and reported an enhancement of factuality by 14.3%. Waldendorf et al. [177] applied contrastive decoding in multilingual machine translation settings. Their approach maximizes the log-likelihood difference between an expert model and a deliberately source-detached amateur model, which significantly enhances translation fidelity and mitigates hallucinations. Furthermore, Sennrich et al. [178] proposed source-contrastive and language-contrastive decoding methods, contrasting the correct input segment with randomly selected or incorrect segments. This approach substantially reduced severe hallucinations (defined by a chrF2 score below 10) by 67–83% and oscillatory hallucinations by 75–92% across various language pairs.
@@
Knowledge Distillation. It is a method whereby a smaller model, the student, learns to replicate the performance of a larger, well-performing model, the teacher, without significant performance loss [179]. Recent studies have shown knowledge distillation to be effective in mitigating hallucinations in LLMs. McDonald et al. [180] employed knowledge distillation with the Mistral LLM, demonstrating substantial improvements in factual accuracy and significant reductions in hallucination rates on the MMLU benchmark. The proposed approach used temperature scaling and intermediate layer matching, which enables a compact student model to emulate a larger teacher model. Therefore, the approach ensures more contextually accurate outputs without compromising computational efficiency. Similarly, Nguyen et al. [181] proposed a smoothed knowledge distillation method, where soft labels from a teacher model replaced traditional hard labels, to reduce the model’s overconfidence and encourage better factual grounding. This method was evaluated on summarization benchmarks, such as CNN/Daily Mail and XSUM, and it successfully lowered hallucination rates while maintaining robust performance across general natural language processing tasks. Liu et al. [182] designed a multi-task learning (MTL) paradigm to distill self-evaluation capabilities from GPT-3.5-turbo into smaller models like T5-base. This approach uses few-shot CoT prompts and self-assessment outputs to create rationales and pseudo-labels for training. Results across the SVAMP and ANLI datasets demonstrated significant improvements in accuracy. Elaraby et al. [183] introduced HALO, a framework addressing hallucinations in smaller LLMs like BLOOM 7B through HALOCHECK, which is a BlackBox knowledge-free metric for evaluating hallucination severity. HALOCHECK uses entailment-based methods to assess consistency, which outperformed existing metrics in detecting contradictions. To mitigate hallucinations, the authors proposed knowledge injection, fine-tuning the model with domain-specific knowledge, and a teacher-student approach where GPT-4 provided detailed guidance selectively triggered by HALOCHECK. These techniques significantly improved factual consistency and reduced hallucinations, demonstrating HALO’s effectiveness in enhancing weak LLMs’ reliability in domain-specific tasks.
@@
Supervised Fine-Tuning. Although the supervised fine-tuning can produce hallucinated text, it can be an effective hallucination mitigation strategy when specific factors are carefully considered. Incorporating grounded input into the fine-tuning dataset enables the model to prioritize factually accurate content over speculative knowledge [184]. For example, if the model frequently hallucinates in text summarization tasks, fine-tuning with a dataset containing grounded content is beneficial, since it guides the model in aligning generated summaries closely with the original content.
@@
Training the model to associate uncertainty with lower confidence scores is helpful in uncertain situations [185]. For instance, while summarizing a detailed medical article with specialized terms and domain-specific information, the model can be trained to flag areas of uncertainty. As a result, the model may either omit this information from the summary or openly indicate limitations (e.g., ”certain details were not specified”). Similarly, instructional supervision can support this approach by teaching the model that disclaiming uncertain information is preferable to generating incorrect data. This method allows the model to respond with ”I don’t know” when information is unavailable [186].
@@
The hallucination issue may stem from the dependence of the training process on a single dataset, which limits the model’s ability to grasp the true features of the task. Incorporating appropriate auxiliary tasks alongside the primary task during fine-tuning helps reduce the model’s susceptibility to hallucination issues [5]. The process of fine-tuning a model on several tasks concurrently is known as MTL. Learning from multiple tasks simultaneously enables models to gather general information beyond task-specific properties. Consequently, learning performance can improve by sharing information across tasks rather than learning each task separately. For example, in an MTL framework, encompassing abstractive summarization and fact-checking makes the model learn to produce coherent summaries and to validate the factual accuracy of the information it generates. The fact-checking task helps prevent the model from producing erroneous or fabricated statements by explicitly instructing it to assess and verify factual assertions. However, it is essential to carefully select tasks for concurrent learning to avoid the risk of negative transfer, which occurs when learning conflicting features [187].
@@
Prompt retrieval and selection techniques integrated with fine-tuning have further expanded the capabilities of LLMs in mitigating hallucinations [188]. Cheng et al. [188] proposed UPRISE, which is a method to detect and mitigate hallucinations by enhancing the retrieval of relevant prompts that assist the LLM in making more accurate conclusions. The authors fine-tuned a lightweight model to autonomously extract prompts from a pool of prompts based on a zero-shot task input. The model has two encoders: one for processing the task input and the other for processing the prompt. During training, the model maximizes the similarity between the task input and positive prompts while minimizing the similarity to negative prompts through contrastive learning. UPRISE evaluates retrieval scores according to the LLM’s accuracy in anticipating the proper label. Upon fine-tuning the retriever, the fine-tuned retriever is employed to extract the most relevant prompts from the prompt pool for a certain input. The obtained prompts are combined with the task inputs and forwarded to the LLM to produce the final output. The proposed method was evaluated using ChatGPT, and UPRISE outperformed vanilla zero-shot prompting on fact-checking tasks.

### 18 Mar 2026 v3

@@ 7.2 Retrieval-Based Techniques
Retrieving external knowledge for LLMs can enhance their factual grounding, contextual relevance, and overall accuracy. The information contained in the model parameters during pre-training is known as parametric knowledge. This knowledge is static and may not encompass the most recent or domain-specific information. To overcome this limitation, non-parametric knowledge can be incorporated through retrieval-based techniques. In the context of LLMs, knowledge retrieval enables them to dynamically access external, verified sources of data, such as search engines, databases, and specific text corpora. By interacting with these reliable sources, models can consult up-to-date and domain-specific data during generation, which in turn improves the factual accuracy of their outputs [154, 155]. Retrieval-based techniques can be classified into three categories based on the phase of accessing external resources: before, during, and after generation [9]. We used these knowledge retrieval stages to categorize retrieval-based techniques proposed for hallucination mitigation.
@@ 7.2.2 Knowledge Graph
Inference stage. KGs can also mitigate hallucinations in the inference stage by employing KG-augmented retrieval. This allows LLMs to retrieve the appropriate subgraph for answering a query grounded by real-world facts. KAPING [174] retrieves related triples from KGs for zero-shot QA by matching entities in questions. Similarly, StructGPT [175] enhances LLM responses using data from KGs, tables, and databases and employing structured queries for information retrieval. In another study, KG is used to enhance evidence acquisition [176]. CuriousLLM [176] is proposed, which builds and traverses a passage-level KG where nodes represent documents and edges reflect learned semantic connections. The LLM agent generates follow-up questions that guide traversal over this graph to locate missing evidence across documents. The proposed technique actively identifies what information is still missing and continues retrieval until sufficient support is found, or terminates early when evidence is complete.
For complex reasoning tasks, IRCoT [177] integrates KGs with CoT to iteratively guide retrieval and reasoning for multi-step questions. Likewise, RoG [178] uses KGs to construct reliable reasoning pathways grounded in diverse relationships to improve accuracy. Moreover, PoG [179] retrieves a question-specific subgraph for each query, explores multi-hop reasoning paths between relevant entities, and feeds these structured paths into the LLM as evidence for answer generation. Hallucinations are mitigated by constraining the model to reason over verifiable KG paths, pruning noisy or irrelevant branches, and verifying whether the retrieved paths provide sufficient support before producing an answer. For real-time applications, FactGenius [180] combines LLM-based connection filtering with Levenshtein distance validation, which improved fact verification F1-scores by 12% on the FactKG benchmark through fuzzy relation mining. Moreover, KG-based retrofitting (KGR) [181] incorporates LLMs with KGs to mitigate factual hallucination during the reasoning process. Unlike previous methods that rely solely on user input to query the knowledge graph, KGR retrofits LLM initial draft responses using factual knowledge stored in KGs. The framework leverages LLMs to extract, select, validate, and retrofit factual statements in model-generated responses, enabling an autonomous knowledge-verification and refinement process without additional manual effort.
@@ 7.3 Reasoning-Based Techniques
CoT Reasoning. CoT is an improved version of few-shot prompting [145] that follows a step-by-step approach in natural language to decompose multi-step problems into intermediate steps. This method provides an interpretable view of the model’s reasoning process, which enables better insight into how the model arrives at its conclusions. Zhao et al. [182] proposed a zero-shot CoT reasoning method by incorporating logical thoughts (LoT), which employs symbolic logic to validate and revise each step. Therefore, it addresses and mitigates hallucinations arising from logical flaws directly. Sultan et al. [183] introduced Structured CoT (SCoT), which uses a structured state-machine methodology to systematically oversee content reading, utterance generation, and hallucination mitigation phases. This method improves the model’s faithfulness and reduces hallucinations in content-grounded tasks. Li et al. [184] further adapted the SCoT approach specifically for code generation tasks, introducing programming structures, such as sequential, branching, and looping steps, as intermediate representations. Their approach explicitly guides LLMs to follow structured programming logic, which enhances the logical correctness, clarity, and readability of the generated code. This structured reasoning substantially mitigates hallucinations and improves code accuracy, outperforming traditional CoT prompting by a considerable margin across multiple programming benchmarks. Li et al. [185] improved the causal reasoning capabilities of LLMs by introducing CDCR-SFT. The authors fine-tune LLMs to explicitly construct causal graphs and reason over them before producing answers. By shifting reasoning from surface token patterns to structured causal relationships, the model reduces logically inconsistent hallucinations that arise from flawed reasoning chains.
@@
Self-Consistency. Building on the success of CoT, Wang et al. [186] introduced a self-consistency approach, which replaces the decoding strategy of CoT. Unlike the original CoT method, the authors proposed the concept of generating multiple reasoning chains for a certain topic. The final answer is determined by conducting a majority vote or by selecting the most consistent response from several chains. Li et al. [187] extended self-consistency by proposing DIVERSE. This method generates multiple reasoning pathways through diverse prompts, systematically validates each reasoning step, and utilizes weighted voting according to step consistency. The step-aware approach significantly reduces hallucinations by isolating and rectifying errors at the step level rather than the whole chain. Similarly, Liang et al. [188] proposed an RL from knowledge feedback (RLKF) technique, which integrates self-consistency and internal knowledge state assessments. RLKF enhances the model’s ability to resist factual hallucinations by reinforcing the consistency between the internal knowledge state and external outputs. Xu et al. [189] proposed SaySelf, a method for training LLMs to produce self-reflective rationales conditioned on inconsistencies across multiple sampled reasoning trajectories. SaySelf explicitly prompts the model to provide precise confidence estimates and rationales that detail its knowledge gaps and uncertainty. This method substantially improves confidence calibration and effectively mitigates hallucinations.
@@
Iterative Refinement. CoT has substantially improved the reasoning abilities of LLMs, thus reducing hallucination. Nevertheless, if CoT initiates the sequence with erroneous reasoning, it fails to rectify the reasoning errors or factual inaccuracies that may occur during the reasoning process. This limitation often leads to hallucinations. Inspired by the way people edit their written content, Madaan et al. [146] proposed Self-Refine to enhance the generated output by an LLM through iterative refinement, which depends on domain-specific data, external supervision, and RL. However, these techniques require large annotated datasets that are unavailable for several domains. Self-Refine leverages the concept of LLMs as self-reviewers by using a single LLM that acts as a generator, refiner, and feedback provider, relying solely on prompt designs without the need for fine-tuning. Self-Refine depends on an appropriate LLM and three prompts (for initial response generation, feedback, and refinement). It works by generating an initial output based on an input sequence, providing feedback on the output, then subsequently refining the output in accordance with the self-generated feedback. It alternates between feedback and refining until a specified condition is achieved. Similarly, Shinn et al. [190] proposed Reflexion method that involves finding the subsequent optimal solution in planning using ReAct [191]. The Reflexion uses a long-lasting memory that allows an agent to recognize its own mistakes and autonomously derive insights from its mistakes and iteratively adapt its behavior over time. Instead of using RL, Reflexion leverages an external verbal feedback mechanism. This technique operates by employing an actor that generates an output and an evaluator to evaluate the generated output. The process continues until the evaluator thinks the agent’s output is correct or when a maximum number of trials is reached. In contrast to Self-Refine, Reflexion explicitly integrates external evaluations and memory mechanisms, which enhances learning effectiveness over both short-term and long-term experiences. Paul et al. [192] developed REFINER by applying structured intermediate criticism from a critic model. The critic model critiques single reasoning steps and allows iterative refinement by identifying and correcting inaccuracies throughout the reasoning process. Likewise, Lee et al. [193] introduced the Ask, Assess, and Refine (A2R) methodology, which methodically assesses outputs for factual accuracy and hallucinations through metric-based feedback. A2R employs natural language feedback from metric assessments, which progressively enhances responses according to metrics like accuracy, fluency, and citation precision to reduce hallucination. In another study, Zhang et al. [194] introduced PREFER, a refinement-focused ensemble method, which works by continuously merging and refining prompts through identifying the limitations of previous iterations, enhancing model stability and overall output quality. In addition, Huang et al. [195] demonstrated that LLMs can enhance their reasoning capabilities by generating high-confidence rationale-augmented responses. Their approach integrates unsupervised iterative refinement with self-consistency and feedback-based fine-tuning, which significantly improved reasoning performance. Cheng et al. [196] proposed HaluSearch models generation as a slow, step-by-step reasoning search using tree search (e.g., MCTS). Each intermediate reasoning step is scored by a reward model that estimates hallucination risk, and the model dynamically switches between fast and slow thinking. This prevents early reasoning errors from propagating and promotes more reliable reasoning paths.
@@ 7.4 Model-centric Training and Adaptation
Decoding Strategies. Decoding is the mechanism used to convert encoded representations of the LLMs’ output into comprehensible text. During decoding, the model iteratively selects tokens from its vocabulary, which build contextually relevant and syntactically accurate sentences. However, decoding strategies may contribute to the generation of hallucinated text [76]. Selecting an effective decoding technique can help the model generate output that is more grounded in context and aligned with user expectations. Recent studies have shown promising results in mitigating hallucinations by building upon the on-the-shelf decoding strategies. Chen et al. [203] introduced DoLa to mitigate hallucination in LLMs without the need for fine-tuning or external retrieval mechanisms. DoLa enhances factual accuracy by exploiting differences between mature (higher) and premature (lower) transformer layers. Evaluated on the TruthfulQA benchmark, DoLa attained a 12–17% improvement in the truthfulness scores in various LLMs. Similarly, Shi et al. [204] enhanced the contextual faithfulness of text generation by reducing the influence of prior knowledge during the decoding process. They proposed a contrastive decoding technique that modifies output probabilities to enhance the influence of the given context. The context-aware decoding strategy operates by contrasting the output probabilities of the model when the context is included and excluded from the prompt. This ensures that the model gives more weight to contextually relevant tokens, which reduces reliance on outdated or incorrect prior knowledge. This approach has been evaluated on a summarization task and reported an enhancement of factuality by 14.3%. Waldendorf et al. [205] applied contrastive decoding in multilingual machine translation settings. Their approach maximizes the log-likelihood difference between an expert model and a deliberately source-detached amateur model, which significantly enhances translation fidelity and mitigates hallucinations. Furthermore, Sennrich et al. [206] proposed source-contrastive and language-contrastive decoding methods, contrasting the correct input segment with randomly selected or incorrect segments. This approach substantially reduced severe hallucinations, which is defined by a chrF2 score below 10, by 67–83% and oscillatory hallucinations by 75–92% across various language pairs. In order to reduce the response latency of existing methods, Chang et al. [207] introduced monitoring decoding, which monitors token-by-token generation using a factuality monitor that evaluates partial responses during decoding. When a token or span is predicted to cause hallucination, the model intervenes immediately using a tree-based resampling strategy to revise only the risky tokens instead of regenerating the whole answer. This reduces the number of overconfident hallucinated tokens at their source during generation.
@@
Knowledge Distillation. It is a method by which a smaller model, the student, learns to replicate the performance of a larger, well-performing model, the teacher, without significant performance loss [208]. Recent studies have shown that knowledge distillation is effective in mitigating LLM hallucinations. McDonald et al. [209] used knowledge distillation with the Mistral LLM, demonstrating substantial improvements in factual accuracy and significant reductions in hallucination rates on the MMLU benchmark. The proposed approach used temperature scaling and intermediate layer matching, which enables a compact student model to emulate a larger teacher model. Therefore, the approach ensures more contextually accurate outputs without compromising computational efficiency. Similarly, Nguyen et al. [210] proposed a smoothed knowledge distillation method, where soft labels from a teacher model replaced traditional hard labels, to reduce the model’s overconfidence and encourage better factual grounding. This method was evaluated on summarization benchmarks, such as CNN/Daily Mail and XSUM, and it successfully lowered hallucination rates while maintaining robust performance across general natural language processing tasks. Liu et al. [211] designed a multi-task learning (MTL) paradigm to distill self-evaluation capabilities from the GPT-3.5-turbo into smaller models such as the T5-base. This approach uses few-shot CoT prompts and self-assessment output to create rationales and pseudo-labels for training. The results in the SVAMP and ANLI datasets demonstrated significant improvements in accuracy. Elaraby et al. [212] introduced HALO, a framework addressing hallucinations in smaller LLMs such as BLOOM 7B through HALOCHECK, which is a BlackBox knowledge-free metric to evaluate hallucination severity. HALOCHECK uses entailment-based methods to assess consistency, which outperformed existing metrics in detecting contradictions. To mitigate hallucinations, the authors proposed knowledge injection, fine-tuning the model with domain-specific knowledge, and a teacher-student approach in which GPT-4 provided detailed guidance selectively triggered by HALOCHECK. These techniques significantly improved factual consistency and reduced hallucinations, demonstrating HALO’s effectiveness in enhancing weak LLMs’ reliability in domain-specific tasks.
@@
Supervised Fine-Tuning. Although supervised fine-tuning can produce hallucinated text, it can be an effective hallucination mitigation strategy when specific factors are carefully considered. Incorporating grounded input into the fine-tuning dataset enables the model to prioritize accurate factual content over speculative knowledge [213]. For example, if the model frequently hallucinates in text summarization tasks, fine-tuning with a dataset containing grounded content is beneficial, since it guides the model in aligning generated summaries closely with the original content.
@@
Uncertainty-aware Learning. Training the model to associate uncertainty with lower confidence scores is helpful in uncertain situations [214]. For instance, while summarizing a detailed medical article with specialized terms and domain-specific information, the model can be trained to flag areas of uncertainty. As a result, the model may either omit this information from the summary or openly indicate limitations (e.g., ”certain details were not specified”). Similarly, instructional supervision can support this approach by teaching the model that disclaiming uncertain information is preferable to generating incorrect data. This method allows the model to respond with ”I don’t know” when information is unavailable [215, 216]. Dey et al. [217] introduced uncertainty-aware fusion (UAF), which reduces hallucination using an ensemble of multiple LLMs combined through uncertainty estimation. Each model provides both an answer and a confidence signal about the factual correctness. A fusion module then selects or combines responses from models that are both accurate and self-aware. The proposed approach enhanced factuality while also maintaining efficiency.
@@
Multi-task Learning. The hallucination issue may stem from the dependence of the training process on a single dataset, which limits the model’s ability to grasp the true features of the task. Incorporating appropriate auxiliary tasks alongside the primary task during fine-tuning helps reduce the model’s susceptibility to hallucination issues [5]. The process of fine-tuning a model on several tasks concurrently is known as MTL. Learning multiple tasks simultaneously enables models to gather general information beyond task-specific properties. Consequently, learning performance can improve by sharing information across tasks rather than learning each task separately. For example, in an MTL framework, encompassing abstractive summarization and fact-checking makes the model learn to produce coherent summaries and to validate the factual accuracy of the information it generates. The fact-checking task helps to prevent the model from producing erroneous or fabricated statements by explicitly instructing it to assess and verify factual assertions. However, it is essential to carefully select tasks for concurrent learning to avoid the risk of negative transfer, which occurs when learning conflicting features [218].
@@
Prompt-guided Learning. Prompt retrieval and selection techniques integrated with fine-tuning have further expanded the capabilities of LLMs to mitigate hallucinations [219]. Cheng et al. [219] proposed UPRISE, which is a method of detecting and mitigating hallucinations by enhancing the retrieval of relevant prompts that help the LLM in making more accurate conclusions. The authors fine-tuned a lightweight model to autonomously extract prompts from a pool of prompts based on a zero-shot task input. The model has two encoders: one for processing the task input and the other for processing the prompt. During training, the model maximizes the similarity between task inputs and positive prompts while minimizing the similarity to negative prompts through contrastive learning. UPRISE evaluates retrieval scores according to the LLM’s accuracy in anticipating the proper label. Upon fine-tuning the retriever, the fine-tuned retriever is employed to extract the most relevant prompts from the prompt pool for a certain input. The prompts obtained are combined with the task inputs and forwarded to the LLM to produce the final output. The proposed method was evaluated using ChatGPT, and UPRISE outperformed vanilla zero-shot prompting on fact-checking tasks.